# ?reference?

### v1

@@ Abstract
Automating the creation of scientific diagrams from academic papers can significantly streamline the development of tutorials, presentations, and posters, thereby saving time and accelerating the process. Current text-to-image models struggle with generating accurate and visually appealing diagrams from long-context inputs. We propose SciDoc2Diagram, a task that extracts relevant information from scientific papers and generates diagrams, along with a benchmarking dataset, SciDoc2DiagramBench. We develop a multi-step pipeline SciDoc2Diagrammer that generates diagrams based on user intentions using intermediate code generation. We observed that initial diagram drafts were often incomplete or unfaithful to the source, leading us to develop SciDoc2Diagrammer-Multi-Aspect-Feedback (MAF), a refinement strategy that significantly enhances factual correctness and visual appeal and outperforms existing models on both automatic and human judgement.
@@ 1 Introduction
Researchers often need to transform detailed studies into compelling diagrams for presentations, going beyond the usual figures in their papers to include summaries, methodologies, and more dynamic data visualizations Chapman et al. (2014). [^1] According to Fu et al. (2021), approximately 39.2% of figures in ACL slides are creatively derived from the original papers, taking forms like flowcharts, summary tables, and plots/charts that visually interpret the paper’s content. Automation of diagram generation from the multimodal components of academic papers—such as text, figures, tables, and plots—could greatly facilitate the creation of accurate and appealing visual aids for scientific tutorials and presentations Mondal et al. (2024); Fu et al. (2021); Sun et al. (2021), which can help researchers and educators to improve productivity and enhancing the overall quality of their presentations. We introduce a new task SciDoc2Diagram where the input comprises of the content of academic papers and user’s intent and the output is a scientific diagram crafted from the content of papers and aligned with user’s intent, where we refer as “Extrapolated-Diagram” (Section 2).
Figure 1: A busy scientist is inclined to automate the creation of diagrams, where our refinement-based approach in SciDocDiagrammarMAF yields a complete and faithful diagram.
@@
Consider a scenario (Figure 1) where a busy scientist wants to create figures from papers within a tight deadline, with the specific goal of generating a flowchart as shown in the figure (A small-scale pilot study conducted with researchers in the community, as described in 9, reveals the need for fast and accurate generation of scientific figures). To achieve this quickly, recent advancements in text-to-image (T2I) generation models like Stable Diffusion Rombach et al. (2022a) and DALL-E Ramesh et al. (2021, 2022a) offer promising solutions. However, these models face limitations in generating factually correct and visually appealing scientific figures. They primarily produce raster graphics at low resolutions, which is not suitable for scientific figures requiring precision and valid text Hsu et al. (2021). Belouadi et al. (2023) introduced a scientific diagram generation method based on only user’s intent. However, creation of informative diagrams using long-content papers as an additional input remains unexplored.
@@
At the heart of the matter, there is a crucial question: How effective are the current state-of-the-art Large Language models (LLMs) or Visual Language Models (VLMs) at generating scientific diagrams? Despite their potential, a significant knowledge gap exists: there is currently no standard benchmark to assess their performance in this area. To address this gap, we introduce the SciDoc2DiagramBench dataset that features 1080 diagrams from 89 *ACL papers, sourced from existing slides in ACL Anthology (Section 3). This dataset supports researchers in automating scientific diagram creation for presentations.
@@
In this work, we evaluate state-of-the-art LLMs Jiang et al. (2023); Touvron et al. (2023); OpenAI (2023); Abdin et al. (2024) to extract relevant information from documents pertinent to the user-specified intent. Specifically, we use a multi-step pipelined method, SciDoc2Diagrammer that first extracts relevant data from academic papers with respect to intent of the user in creating the diagram, generates an intermediate code that renders the final diagram (Section 4.1). However, the first-draft images generated by these models suffer from a range of errors related to hallucination, completeness and readability due to their incapabilities of parsing long-prompts Lee et al. (2024a).
@@
Therefore, we propose a new sequential refinement approach SciDoc2DiagrammarMAF based on the existing SciDoc2Diagrammer model. The new approach is guided by several critic models to provide targeted feedback (multiple aspects such as Completeness, Faithfulness, and Look and Feel) in a step-by-step hierarchical fashion on the output to further enhance the quality of output (Section 4.2). Our multi-aspect feedback refinement approach significantly improves diagram quality, as confirmed by both automatic and human evaluations, demonstrating notable enhancements over existing methods (Section 6). We observe that complex flowcharts and tables, derived from single or multiple documents, exhibit notable enhancements in faithfulness and completeness when refined sequentially. It is particularly advantageous when iteratively synthesizing information from multiple sources to enhance clarity and precision. [^2] Code and data available at https://github.com/Ishani-Mondal/SciDoc2DiagramGeneration.
@@ 2 Task Formulation of SciDoc2Diagram
Our goal is to automatically generate scientific diagrams from academic papers tailored to user specifications. We process the input document $D=\{T,F,\tau,\pi\}$, which includes body text $T$, figures $F$, tables $\tau$, and plots $\pi$, guided by the user intent $I$ (e.g., “create a flowchart showing the methodology” in Figure 1). The output, an Extrapolated-Diagram $\delta$, visually represents content not explicitly illustrated in the paper.
@@
3 SciDoc2DiagramBench - A Diagram Dataset
Creating Extrapolated-Diagram from source papers is an important, yet challenging task. However, no current dataset evaluates state-of-the-art models on this task. To advance research, we have created SciDoc2DiagramBench dataset, repurposing DOC2PPT Fu et al. (2021) (scientific papers and corresponding slides from ACL Anthology) and TutorialBank Fabbri et al. (2018) (tutorials from *ACL conferences) [^3] Data Contamination issues are discussed in Appendix14. Our dataset includes the following three parts and the overall statistics of different diagram types are shown in Table 1 and examples are shown in Appendix 15.
SciDoc2DiagramBench-Gold
Our initial goal was to explore what are the various types of Extrapolated-Diagram usually created by humans while making presentations. So, we hired an expert annotator from Upwork with more than 5 years of authoring NLP-related papers and creating presentations. The annotator was asked to perform two tasks, first to identify slides containing Extrapolated-Diagram (excludes slides that used external information not found in the paper), second to manually write an intent for the corresponding diagram creation (13, 12). Through this annotation process, we collect 265 gold-standard Extrapolated-Diagram from 34 papers where each instance comprises of a tuple of <papername, intent of diagram, gold diagram>. The first three authors manually checked the intents, and only instances that were unanimously deemed reasonable were retained. After analyzing these diagrams, we observe that 32% diagrams were flowcharts illustrating methodology (Extrapolated-Flowchart), 33% comprised of results/plots (Extrapolated-Results), 25% comprised of summary of related works/contributions (Extrapolated-Summary) and 10% was related to clarifying existing architecture (Extrapolated-Architecture).
@@ 3 SciDoc2DiagramBench - A Diagram Dataset
Using insights from SciDoc2DiagramBench-Gold on how presenters transform paper content into diagrams, we used GPT-4 to generate new diagram intents for this subset of data (Table 14). We also provided guidelines for creating Extrapolated Diagrams that blend text and visuals from academic papers. Each intent, reviewed by the first author, aims to guide the creation of practical diagrams. Each instance includes <paperName, intent> but lacks a corresponding human-created gold diagram.
@@ 4 Methodology
Our proposed SciDoc2Diagram task comprises of generating initial first-draft generation of diagrams from documents using base models (Section 4.1), followed by how the diagrams get iteratively refined based on critic modules (Section 4.2).
@@ 4.1 SciDoc2Diagrammer
We introduce SciDoc2Diagrammer, a two-stage pipeline for generating “Extrapolated-Diagrams” from paper content, where we disentangle the diagram planning and generation stages (Upper portion of Figure 16). In the first Diagram Planning stage, a planner VLM takes the PDF (decomposed into sections and figures/tables) and user’s intent and generates a diagram plan. Inspired by the success of extracting information using question-generation and retrieving answers from relevant snippets as done by Krithara et al. (2023); Lee et al. (2023); Qiu et al. (2018); Auer et al. (2023), we also synthesize diagram plan by first generating clarifying questions based on the user’s intent, and then obtain answers of those questions from relevant portions of the paper to come up with question-answer pairs, which we refer to as diagram plan. In the second stage, inspired by the success of LLMs in generating code from natural language instructions Ugare et al. (2024); Maddigan and Susnjak (2023), we use LLMs to take the diagram plan and user intent as inputs to generate an intermediate code and execute the code to finally render the diagram.
@@
Given an input document $D=\{T,F,\tau,\pi\}$ consisting of body text $T$, figures $F$, tables $\tau$, and plots $\pi$, and user intent $I$, the diagram planning phase includes the following sub-steps: [1] Intent Classification: We use an LLM to classify if the intent of the user is related to results or not. For example in Figure 1, the intent is not related to results, but the one in Figure 16 is related to results. [2] Question Generator: Using the intent, we ask LLMs to generate clarification questions to create the diagram conditioned on intent. For example, for the intent in Figure 16, we obtain a set of clarification questions such as “What are the names of the three best performing baselines that you’d like to compare with NASH?”, “Could you provide the precision values for NASH and the three baseline models on the Reuter Dataset for 128 bits?”, “What are the descriptions of each of these models?”. [3] Retrieve Information from paper and Content Simplication: We retrieve relevant passages from the paper based on each of the generated questions. For instance, in Figure 16, a plot and two paragraphs from Experimental Section of the source paper are retrieved to answer the above questions. If any result type figure is relevant to the question, we first extract data from the figures and convert them into tables for simplification of content retrieval as motivated by prior works Liu et al. (2023a); Masry et al. (2023). [4] Diagram Plan Synthesis: After generating questions and extracting relevant content from paper, we extract answers of those questions. Then we formulate a dictionary of QA pairs/diagram plan $Q=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{n},a_{n})\}$, where $q_{i}$ are the questions and $a_{i}$ are the answers.
@@
In this phase, we leverage the user intent $I$ and QA pairs $Q$ from the previous step to generate an intermediate code $\Theta$ that outlines the diagram construction process. This intermediate code is represented as Python code, as shown in Fig. 16. Finally, we execute the intermediate code $\Theta$ to render the final “Extrapolated-diagram” $\delta$, ensuring that the resulting diagram is both visually appealing and factually accurate, as exemplified by the bar chart at the top-right corner of Figure 16.
Figure 2: The figure (an example from SciDoc2DiagramBench) depicts SciDoc2Diagrammer-MAF, which refines scientific diagrams based on user intent and source document content. Initially, the base model, SciDoc2Diagrammer, creates a diagram (Step 2), which is then refined through three critic modules: Completeness ($C_{critic}$), Faithfulness ($F_{critic}$), and Layout ($L_{critic}$). These modules assess and provide feedback on necessary components, data accuracy, and visual design, respectively. The diagram is repeatedly refined, as shown on the left side of the figure (illustrating SumMAF), where feedback from all critics is integrated at Step 4’. The refinement continues, evaluated at Step 5’, until the maximum iterations are reached or the diagram meets the specified standards.
@@ 4.2 SciDoc2Diagrammer-MAF
Motivation: To gain a qualitative understanding of the diagrams generated from SciDoc2Diagrammer, we randomly sample 50 diagrams from the test set for manual inspection. We found certain frequently occurring errors, like missing components, misleading information, and logical inconsistencies in diagrams like flowcharts. These errors are categorized and detailed in Table 10. We classify these issues into three broad categories: Completeness (essential components are missing or incomplete, affecting the diagram’s informativeness), Faithfulness (misleading information that do not accurately reflect the source data), and Layout (poor visual alignment hindering readability and comprehension). Besides, we also noticed that majority of errors propagate during the transition from our diagram planning to rendering stage; this was due to the inability of LLMs to handle long-context prompts Nathani et al. (2023); Gani et al. (2024).
Inspired by the self-refinement strategies by Madaan et al. (2023), we wanted to study whether the quality of our diagrams can be enhanced by utilizing generic feedback for a wide range of errors or by applying targeted feedback for specific error categories. For exploring the second strategy, we introduce an iterative refinement framework, depicted in Figure 16 (Lower part), designed to enhance quality of diagram generation by specifically rectifying above-mentioned error types.
@@
As illustrated in Figure 2, our MAF framework comprises of the following components: a base model (SciDoc2Diagrammer) that generates an initial diagram ($O$) using both intent and source document (Step 1 in  2). It passes through three specialized feedback critic evaluator modules, each focusing on a critical aspect of diagram quality such as completeness ($C_{critic}$), faithfulness ($F_{critic}$), and layout ($L_{critic}$) on $D$, a Refiner module $R$ that generates refined solution $O\textquoteright$ ’ based on $O$ and feedback.
@@
The completeness critic $C_{critic}$ (Algorithm 1) ensures that a diagram comprehensively represents all critical aspects of the associated content, providing feedback for enhancements based on identified gaps. It is evaluated based on on user’s intent $I$, diagram and source document (Step 3). The process begins by decomposing the diagram’s intent and generating a set of questions $Q$ from the user intent $I$ to determine how well various elements are represented in the diagram. It then involves extracting answers from the diagram, scoring each response for adequacy, calculating an average score as the completeness score $C_{\text{Score}}$, and providing textual feedback ($C_{\text{Feedback}}$) to suggest improvements (Step 4).
The Faithfulness critic $F_{critic}$ (Algorithm 2) evaluates how accurately a diagram represents information from its corresponding document source. It evaluates the accuracy of a diagram by generating questions, retrieving answers from both the document and the diagram, and scoring their consistency to provide an average score and feedback for improvement. The layout critic $L_{critic}$ (Algorithm 4) evaluates how good is the look and feel of generated diagram, and provides score and feedback for improvement. We use two refinement techniques:
@@
Inspired by Nathani et al. (2023), we adopt the summarization-based multi-aspect refinement strategy (Algorithm 4) that iteratively enhances a diagram by integrating feedback from multiple critic modules. Our method is different from theirs in the way that we employ our critic models that act as human proxies, revisiting the PDF source and user intent repeatedly to validate and refine the diagram $C_{critic}$(Algorithm 1) and $F_{critic}$(Algorithm  2), mimicking the iterative human process of cross-verifying information through continuous questioning and answering.
In this algorithm, the first draft is sent to $C_{critic}$ (Algorithm 1) which checks if all necessary components are present, (for instance in Figure 2, it checks the completeness of creating a flowchart). At the same time, the faithfulness evaluator $F_{critic}$ (Algorithm  2) ensures the accuracy of the diagram by comparing it with the source PDF, whether or not all the information in the source figure is true. This is achieved by generating questions (FQ1, FQ2 in Figure 16) based on the diagram and obtaining answers from both the PDF and the diagram. Also, the layout evaluator checks the visual arrangement, ensuring good alignment and spacing (Figure 16). All the critic models provide feedback on a single output (Step 3’) and sent to the refiner. The refiner model makes adjustments by summarizing the suggested modifications from all the critics and generates a new diagram (Step 4’). This process is repeated (Steps 3’, 4’, 5’ in Figure 2) until the evaluator models are satisfied or we reach maximum iterations.
@@
Algorithm 3 is designed to methodically enhance a diagram in a step-by-step hierarchical fashion ensuring that each critic is satisfied (See Flowchart enhancement in Figure 1). The intuition is that, complex tables or flowcharts benefit from being accurate and detailed in their depiction of processes or relationships, where sequential refinement allows for the gradual addition of details, ensuring that each element is correctly represented and connected. As stated in Algorithm 3, the diagram is feed into different critic modules one by one, refine the output until each of them are satisfied.
In Figure 2, $C_{critic}$ first provides feedback (detail composition operators or hierarchical representation, explicit modelling and long range dependency step inclusion) on the diagram in Step 4, and the refiner takes the initial figure and refines based on feedback in Step 5, and passed onto $C_{critic}$ to further evaluate and repeat the same steps (4, 5, 6) till maximum iteration is used or the $C_{critic}$ provides a score greater than 4.5 (as stated in Algorithm 1. The same process is repeated by the $F_{critic}$ and $L_{critic}$ modules and finally generates the output after Step 12.

### v2

@@ Abstract
Automating the creation of scientific diagrams from academic papers can streamline the development of tutorials, presentations, and posters. Current text-to-image models struggle with generating accurate and visually appealing diagrams from long-context inputs. We propose SciDoc2Diagram, a task that extracts relevant information from scientific papers and generates diagrams, along with a benchmarking dataset, SciDoc2DiagramBench. We develop a multi-step pipeline SciDoc2Diagrammer that generates diagrams based on user intent using intermediate code generation. We observed that initial diagram drafts are often incomplete or unfaithful to the source, requiring SciDoc2Diagrammer-Multi-Aspect-Feedback (MAF), a refinement strategy that significantly enhances factual correctness and visual appeal and outperforms existing models on both automatic and human judgement.
@@ 1 Introduction
Researchers often need to transform detailed studies into compelling diagrams for presentations, going beyond the usual figures in their papers to include summaries, methodologies, and dynamic data visualizations Chapman et al. (2014). [^1] According to Fu et al. (2021), approximately 39.2% of figures in ACL slides are creatively derived from the original papers, taking forms like flowcharts, summary tables, and plots/charts that visually interpret the paper’s content. Automation of diagram generation from the multimodal components of academic papers—such as text, figures, tables, and plots—could enable the creation of accurate and appealing visual aids for scientific tutorials and presentations Mondal et al. (2024); Fu et al. (2021); Sun et al. (2021), which can help researchers and educators to improve productivity and enhancing the overall quality of their presentations. We introduce a new task SciDoc2Diagram where the input comprises of the content of academic papers and user’s intent and the output is a scientific diagram crafted from the content of papers and aligned with user’s intent, which we call “Extrapolated-Diagram”.
Figure 1: An example of Diagram Generation and Refinement using SciDoc2Diagrammer-MAF with input as a paper and user-defined intent.
@@
Consider a scenario where a busy scientist wants to create figures from papers within a tight deadline, with the specific goal of generating a bar chart as shown in the Figure 1. To achieve this quickly, recent advancements in text-to-image (T2I) generation models like Stable Diffusion Rombach et al. (2022a) and DALL-E Ramesh et al. (2021, 2022a) offer promising solutions. However, these models face limitations in generating factually correct and visually appealing scientific figures. They primarily produce raster graphics at low resolutions, which is not suitable for scientific figures requiring precision and valid text Hsu et al. (2021). Belouadi et al. (2023) introduced a scientific diagram generation method based on only user’s intent. However, creation of informative diagrams using long-content papers as an additional input remains unexplored.
@@
How effective are the current state-of-the-art Large Language models (LLMs) or Visual Language Models (VLMs) at generating scientific diagrams? Despite their potential, a significant knowledge gap exists: there is currently no standard benchmark to assess their performance in this area. To address this gap, we introduce the SciDoc2DiagramBench dataset that features 1080 diagrams from 89 *ACL papers, sourced from existing slides in ACL Anthology (Section 3). This dataset supports researchers in automating scientific diagram creation for presentations.
@@
We evaluate state-of-the-art LLMs Jiang et al. (2023); Touvron et al. (2023); OpenAI (2023); Abdin et al. (2024) to extract relevant information from documents pertinent to the user-specified intent. Specifically, we use a multi-step pipelined method, SciDoc2Diagrammer that first extracts relevant data from academic papers corresponding to intent of the user in creating the diagram, generates an intermediate code that renders the final diagram (Section 4.1) as shown in Figure 1. However, the first-draft images generated by these models suffer from a range of errors: hallucination, incompleteness and unreadability due to their incapabilities of parsing long-prompts Lee et al. (2024a).
@@
Therefore, we propose a sequential refinement approach SciDoc2Diagrammer-MAF based on the intial draft generated by SciDoc2Diagrammer. The new approach is guided by several critic models to provide targeted feedback (multiple aspects such as Completeness, Faithfulness, and Look and Feel) in a step-by-step hierarchical manner on the output to further enhance the quality of output (Section 4.2). Our multi-aspect feedback refinement approach significantly improves diagram quality, as confirmed by both automatic and human evaluations, demonstrating notable enhancements over existing methods (Section 6). Complex flowcharts and tables derived from single or multiple documents show notable enhancements in faithfulness and completeness when refined sequentially. It is particularly advantageous when iteratively synthesizing information from multiple sources to enhance clarity and precision. [^2] Code and data available at https://github.com/Ishani-Mondal/SciDoc2DiagramGeneration.
@@ 2 Task Formulation of SciDoc2Diagram
Our goal is to automatically generate scientific diagrams from academic papers tailored to user specifications. We process the input document $D=\{T,F,\tau,\pi\}$, which includes body text $T$, figures $F$, tables $\tau$, and plots $\pi$, guided by a user intent $I$ (e.g., “Create a bar chart showing the comparison…” in Figure 18). The output, an Extrapolated-Diagram $\delta$, visually represents content not explicitly illustrated in the paper.
@@
3 SciDoc2DiagramBench: A New Dataset
Creating Extrapolated-Diagram from source papers is an important yet challenging task. However, no current dataset evaluates state-of-the-art models on this task. To advance research, we have created SciDoc2DiagramBench dataset, repurposing DOC2PPT Fu et al. (2021) (scientific papers and corresponding slides from ACL Anthology) and TutorialBank Fabbri et al. (2018) (tutorials from *ACL conferences) [^3] Data Contamination issues are discussed in Appendix 14. Our dataset includes the following three parts and the overall statistics of different diagram types are in Table 1 with examples in Appendix 15.
SciDoc2DiagramBench-Gold.
Our initial goal was to explore what are the types of Extrapolated-Diagram usually created by humans while making presentations. We hired an expert annotator from Upwork with more than five years of authoring NLP-related papers and creating presentations. The annotator had two tasks, first to identify slides containing Extrapolated-Diagram (excludes slides that used external information not found in the paper), second to manually write an intent for the corresponding diagram creation (13, 12). Through this annotation process, we collect 265 gold-standard Extrapolated-Diagram from 34 papers where each instance comprises of a tuple of <papername, intent of diagram, gold diagram>. The first three authors manually checked the intents, and only instances that were unanimously deemed reasonable were retained. After analyzing these diagrams, 32% were flowcharts illustrating methodology (Extrapolated-Flowchart), 33% were results/plots (Extrapolated-Results), 25% were summary of related works/contributions (Extrapolated-Summary) and 10% were related to clarifying existing architecture (Extrapolated-Architecture).
@@ 3 SciDoc2DiagramBench: A New Dataset
Using insights from SciDoc2DiagramBench-Gold on how presenters transform paper content into diagrams, we used GPT-4 to generate new diagram intents for this subset of data (Table 14). We also provide guidelines for creating Extrapolated-Diagram that blend text and visuals from academic papers. Each intent, reviewed by the first author, aims to guide the creation of practical diagrams. Each instance includes <paperName, intent> but lacks a corresponding human-created gold diagram.
@@
Figure 2: Outlines the procedure for generating diagrams from academic papers based on specific user intents, followed by refinement of each critic models.
@@ 4 Methodology
The SciDoc2Diagram task comprises of generating initial first-draft generation of diagrams from documents using base models (Section 4.1), followed by refinement of diagrams using critic modules (Section 4.2).
@@ 4.1 SciDoc2Diagrammer
We introduce SciDoc2Diagrammer, a two-stage pipeline for generating “Extrapolated-Diagrams” from paper content (Upper portion of Figure 2). In the first Diagram Planning stage, a planner VLM takes the PDF (decomposed into sections and figures/tables) and user’s intent and generates a diagram plan. Inspired by the success of extracting information using question-generation and retrieving answers from relevant snippets as done by Krithara et al. (2023); Lee et al. (2023); Qiu et al. (2018); Auer et al. (2023), we also synthesize diagram plan by first generating clarifying questions based on the user’s intent, and then obtain answers of those questions from relevant portions of the paper to come up with question-answer pairs, which we refer to as diagram plan. In the second stage, inspired by the success of LLMs in generating code from natural language instructions Ugare et al. (2024); Maddigan and Susnjak (2023), we use LLMs to take the diagram plan and user intent as inputs to generate an intermediate code and execute the code to finally render the diagram.
@@
Given an input document $D=\{T,F,\tau,\pi\}$ consisting of body text $T$, figures $F$, tables $\tau$, and plots $\pi$, and user intent $I$, the diagram planning phase includes the following sub-steps: [1] Intent Classification: We use an LLM (Few-shot with 4 exemplars) to classify if the intent of the user is related to results or not. For example in Figure 3, the intent is not related to results, but the one in Figure 2 is related to results. [2] Question Generator: Using the intent, we ask LLMs to generate clarification questions to create the diagram conditioned on intent. For example, for the intent in Figure 2, we obtain a set of clarification questions. [3] Retrieve Information: We retrieve relevant passages from the paper based on each of questions Lewis et al. (2020). In Figure 2, a plot and two paragraphs from Experimental Section of the source paper are retrieved to answer the generated questions. If any result type figure is relevant to the question, we first extract data from the figures and convert them into tables for simplification of content retrieval as motivated by prior works Liu et al. (2023a); Masry et al. (2023). [4] Diagram Plan Synthesis: After generating questions and extracting relevant content from paper, we extract answers of those questions. Then we formulate a dictionary of QA pairs/diagram plan $Q=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{n},a_{n})\}$, where $q_{i}$ are the questions and $a_{i}$ are the answers.
@@
In this phase, we use the user intent $I$ and QA pairs $Q$ from the previous step to generate an intermediate code $\Theta$ that constructs diagram. This intermediate code is represented as Python code, as shown in Figure 2. Finally, we execute the intermediate code $\Theta$ to render “Extrapolated-diagram” $\delta$, as exemplified by the bar chart at the top-right corner of Figure 2.
Figure 3: The figure (an example from SciDoc2DiagramBench) depicts SciDoc2Diagrammer-MAF, which refines diagrams based on user intent and source document. Initially, SciDoc2Diagrammer creates a diagram (Step 2), which is refined through three critic modules which assess and provide feedback on necessary components, data accuracy, and visual design. The diagram is repeatedly refined, as shown on the left side of the figure (illustrating SumMAF), where feedback from all critics is integrated at Step 4’. The refinement continues, evaluated at Step 5’, until the maximum iterations are reached or the diagram meets the specified standards.
@@ 4.2 SciDoc2Diagrammer-MAF
Motivation: To gain a qualitative understanding of the diagrams generated from SciDoc2Diagrammer, we randomly sample 50 diagrams from the test set for manual inspection. There are some frequently occurring errors (Table 10) and we classify them into three broad categories: Completeness (essential components are missing or incomplete, affecting the diagram’s informativeness), Faithfulness (misleading information that do not accurately reflect the source data), and Layout (poor visual alignment hindering readability and comprehension). Besides, we also noticed that most of the errors propagate during the transition from our diagram planning to rendering stage; this was due to the inability of LLMs to handle long-context prompts Nathani et al. (2023); Gani et al. (2024).
Inspired by the self-refinement strategies by Madaan et al. (2023), we wanted to study whether the quality of our diagrams can be enhanced by using generic feedback for a wide range of errors or by applying targeted feedback for specific error categories. For exploring the second strategy, we introduce an iterative refinement framework, depicted in Figure 2 (Lower part), designed to enhance quality of diagram generation by specifically rectifying above-mentioned error types.
@@
As illustrated in Figure 3, our MAF framework comprises of the following components: a base model (SciDoc2Diagrammer) that generates an initial diagram ($O$) using both intent and source document (Step 1 in  3). It passes through three specialized feedback critic evaluator modules, each focusing on a critical aspect of diagram quality such as completeness ($C_{critic}$), faithfulness ($F_{critic}$), and layout ($L_{critic}$) on $D$, a Refiner module $R$ that generates refined diagram $O\textquoteright$ ’ based on $O$ and feedback.
@@
The completeness critic $C_{critic}$ (Algorithm 1) decomposes the diagram’s intent $I$ and generates a set of questions $Q$ from the user intent $I$ to determine how well various elements are represented in the diagram (Step 3). It then involves extracting answers from the diagram, scoring each response for adequacy, calculating an average score as the completeness score $C_{\text{Score}}$, and providing textual feedback ($C_{\text{Feedback}}$) to suggest improvements (Step 4).
The Faithfulness critic $F_{critic}$ (Algorithm 2) evaluates how accurately a diagram represents information from its document source. It evaluates the accuracy of a diagram by generating questions, retrieving answers from both the document and the diagram, and scoring their consistency to provide an average score and feedback for improvement. The layout critic $L_{critic}$ (Algorithm 4) evaluates how good is the look and feel of generated diagram, and provides score and feedback for improvement. We use two refinement techniques:
@@
We adopt summarization-based multi-aspect refinement strategy from Nathani et al. (2023) for diagram refinement. However, in contrast to their method, we use our critic models that act as human proxies (Algorithm 4), and revisit PDF source and user intent repeatedly to validate and refine the diagram $C_{critic}$(Algorithm 1) and $F_{critic}$(Algorithm  2), mimicking the iterative human process of cross-verifying information through continuous questioning and answering.
The first draft is sent to $C_{critic}$ (Algorithm 1) which checks if all necessary components are present (In Figure 3, it checks the completeness of creating a flowchart). At the same time, the faithfulness evaluator $F_{critic}$ (Algorithm 2) ensures the accuracy of diagram by comparing it with source PDF, whether or not all the information in the source figure is true, by generating questions (FQ1, FQ2 in Figure 2) based on the diagram and obtaining answers from both PDF and diagram. Also, the layout evaluator checks the visual arrangement, ensuring good alignment and spacing (Figure 2). All the critic models provide feedback on a single output (Step 3’) and sent to the refiner. The refiner model makes adjustments by summarizing the suggested modifications from all critics and generates a new diagram (Step 4’). This is repeated (Steps 3’, 4’, 5’ in Figure 3) until the evaluator models are satisfied or we reach maximum iterations.
@@
Algorithm 3 enhances a diagram in a step-by-step hierarchical manner ensuring that each critic is satisfied (See Figure 18). The intuition is that, complex tables or flowcharts benefit from being accurate and detailed in their depiction of processes or relationships, where sequential refinement allows for gradual addition of details, ensuring that each element is correctly represented. Here the diagram is feed into different critic modules one by one, refine the output until each of them are satisfied.
In Figure 3, $C_{critic}$ first provides feedback (detail composition operators or hierarchical representation, explicit modelling and long range dependency step inclusion) on the diagram in Step 4, and the refiner takes the initial figure and refines based on feedback in Step 5, and passed onto $C_{critic}$ to further evaluate and repeat the same steps (4, 5, 6) till maximum iteration is used or the $C_{critic}$ provides a score greater than 4.5 (as stated in Algorithm 1. The same process is repeated by $F_{critic}$ and $L_{critic}$ modules and generates output after Step 12.